Find a rational zero of the polynomial function and use it to find all the zeros of the function.
f(x)=2x^3-x^2-12x+6

Answer by stanbon(75887)

Find a rational zero of the polynomial function and use it to find all the zeros of the function.
f(x)=2x^3-x^2-12x+6
-------------------------
Using synthetic division:
(1/2)....2....-1....-12....6
..........2.....0......-12..|..0
Zero at x=1/2
--------------
Examine 2x^2-12 = 0
2x^2 = 12
x^2 = 6
Zeroes at x = sqrt(6) and at x = -sqrt(6)
=============================
Chees,
Stan H.